OpenStudy (anonymous):

Trevor solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x − 10 + 4x = 10 5x − 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5 He should have substituted 5 + 2x He combined like terms incorrectly, it should have been 4x instead of 5x He subtracted 10 from the right side instead of adding 10 to the right side He made no mistake

OpenStudy (anonymous):

he subtracted 10 instead of added it

OpenStudy (anonymous):

shold have had 5 x -10 + 10 = 10 + 10 to give 5x = 20

OpenStudy (lalaly):

y = 5 − 2x x − 2(5 − 2x) = 10 x − 10 + 4x = 10 5x − 10 = 10......(u should add 10 to both sides) 5x =10+10 5x=20 x=4